OpenStudy (anonymous):

a chemist has one solution that is 25% acid and a second that is 50% acid. How many liters of each chould be mixed of each to get a 10L of a solution that is 40% acid? WHAT ARE THE STEPS TO BREAKING THIS DOWN

OpenStudy (anonymous):

well, assume he has x L of the first solution, so 25% of x L is acid therefore he has 0.25x L of acid. assume he has y L of the second solution, so he has 50% of y L of acid therefore he has 0.50y L of acid. he wants to mix the two amounts to get x + y L of solution, but he wants that solution to be 40% acid, so he wants to get 0.40(x + y) L of acid 0.25x + 0.50y = 0.40(x + y) hence 0.25x + 0.50y = 0.40x + 0.40y hence 0.15x - 0.10y = 0 therefore 15x - 10y = 0 therefore 3x - 2y = 0 But we also know that x + y = 10 multiply equation 2 by 3: 3x + 3y = 30 now subtract the first from the second: 3x + 3y - 3x + 2y = 30 hence 5y = 30 hence y = 6 thus x = 4 so he should mix 4 L of the 25% solution and 6 L of the 50% solution